### 17 Jan 2024 v1

# Universal contributions to charge fluctuations in spin chains at finite temperature

Abstract
At finite temperature, conserved charges undergo thermal fluctuations in a quantum many-body system in the grand canonical ensemble. The full structure of the fluctuations of the total U(1) charge $Q$ can be succinctly captured by the generating function $G(\theta)=\langle\,\mathrm{e}^{\mathrm{i}\theta Q}\rangle$. For a 1D translation-invariant spin chain, in the thermodynamic limit the magnitude $\left\lvert G(\theta)\right\rvert$ scales with the system size $L$ as $\ln\left\lvert G(\theta)\right\rvert=-\alpha(\theta)L+\gamma(\theta)$, where $\gamma(\theta)$ is the scale-invariant contribution and may encode universal information about the underlying system. In this work we investigate the behavior and physical meaning of $\gamma(\theta)$ when the system is periodic. We find that $\gamma(\theta)$ only takes non-zero values at isolated points of $\theta$, which is $\theta=\pi$ for all our examples. In two exemplary lattice systems we show that $\gamma(\pi)$ takes quantized values when the U(1) symmetry exhibits a specific type of ’t Hooft anomaly with other symmetries. In other cases, we investigate how $\gamma(\theta)$ depends on microscopic conditions (such as the filling factor) in field theory and exactly solvable lattice models.
I Introduction
At finite temperature, the state of a quantum many-body system with $\mathrm{U}(1)$ symmetry is conveniently described by the thermal density matrix $\rho=\frac{1}{Z}\,\mathrm{e}^{-\beta(H-\mu Q)}$ in the grand canonical ensemble, where $\beta$ is the inverse temperature and $\mu$ is the chemical potential. It is well-known that even though the total charge $Q$ is not conserved in the ensemble, it has a well-defined average with fluctuations suppressed in the thermodynamic limit. Higher moments of $Q$ are also suppressed as the distribution approaches a Gaussian one. Our main objective in this work is to show that the charge fluctuations may contain universal information about the system, especially when the $\mathrm{U}(1)$ charge-conservation symmetry exhibits certain types of ’t Hooft anomaly (usually a mixed anomaly with other global symmetries). To extract this information, it is most useful to consider the following generating function (also known as the full counting statistics Belzig and Nazarov (2001); Levitov and Reznikov (2004); Song et al. (2012, 2010))
$G(\theta)=\langle\,\mathrm{e}^{\mathrm{i}\theta Q}\rangle=\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\rho.$ (1)
In (1+1)d, for a system of size $L$, it is expected that $G(\theta)$ takes the following form
$\ln\left\lvert G(\theta)\right\rvert=-\alpha(\theta)L+\gamma(\theta)+\cdots,$ (2)
for large $L$, and the phase factor is defined as
$\omega(\theta)=\frac{G(\theta)}{\left\lvert G(\theta)\right\rvert}.$ (3)
In general, the value of $\alpha(\theta)$ is sensitive to microscopic details and is thus non-universal. On the other hand, $\gamma(\theta)$ and $\omega(\theta)$ are expected to encode universal information about many-body states.
Another motivation to study $G(\theta)$ comes from its connection with the disorder parameter in the (2+1)d ground state of a gapped Hamiltonian Wang et al. (2021); Wu et al. (2021); Chen et al. (2022); Cuomo et al. (2022). Suppose $A$ is a subregion of the (2+1)d system, and denote by $Q_{A}$ the total charge inside $A$. We define the disorder parameter as $\langle\,\mathrm{e}^{\mathrm{i}\theta Q_{A}}\rangle$, where the expectation value now is taken with respect to the ground state. When the ground state is gapped and preserves the $\mathrm{U}(1)$ symmetry, one expects for a region $A$ of size $L_{A}$
$\ln\left\lvert\langle\,\mathrm{e}^{\mathrm{i}\theta Q_{A}}\rangle\right\rvert=-\alpha_{1}(\theta)L_{A}+\gamma(\theta)+\mathcal{O}(1/L_{A}).$ (4)
To make connections to the (1+1)d discussion more explicit, we notice that in many cases the reduced density matrix can be well-approximated by a thermal state of a (quasi-)local Hamiltonian acting on degrees of freedom localized at the boundary of $A$, which is effectively a (1+1)d system. If the (2+1)d bulk is a nontrivial symmetry-protected topological (SPT) state, the symmetries should act anomalously in the effective boundary theory. Thus the computation of $\langle\,\mathrm{e}^{\mathrm{i}\theta Q_{A}}\rangle$ is reduced to $G(\theta)$ in an effective (1+1)d system.
In this work, we study the behavior and the physical interpretations of $\gamma(\theta)$ and $\omega(\theta)$. A key question to address is to what extent the values of $\gamma(\theta)$ are truly universal, i.e. unaffected by small changes to the Hamiltonian. Additionally, we explore the physical significance of these universal, quantized values.
First, we present a general computation of $G(\theta)$ when the (1+1)d system can be described by a conformal field theory (CFT), possibly with a topological defect. In the absence of such a defect, we find that $\gamma(\theta)$ effectively counts the degeneracy of the defect operator of the corresponding $\mathrm{U}(1)$ symmetry transformation. It only takes non-zero values at isolated values of $\theta$, as some of the degeneracies are enforced by the mixed anomaly between $\mathrm{U}(1)$ and other global symmetries.
We then compute $\gamma(\theta)$ and $\omega(\theta)$ in the presence of a $\mathrm{U}(1)$ topological defect for a $c=1$ free boson CFT. It turns out that both $\gamma(\theta)$ and $\omega(\theta)$ are now sensitive to the $\mathrm{U}(1)$ defect, which can be understood as a consequence of the ’t Hooft anomaly of the $\mathrm{U}(1)\times\mathrm{U}(1)$ symmetry of the free boson CFT.
Next we turn to $\gamma(\theta)$ and $\omega(\theta)$ in lattice models, to understand how $\gamma(\theta)$ and $\omega(\theta)$ behave at high temperature beyond the CFT description. We also address the question as to whether the value of $\gamma(\theta)$ is quantized or not. Our main finding is that in the presence of certain types of mixed anomaly between U(1) and other global symmetry, $\gamma(\pi)$ takes quantized values for a general thermal state of any symmetric local Hamiltonian. We establish this result for two types of systems:
  1. 1.
    A translation-invariant spin-1/2 chain with on-site O(2) symmetry, where the O(2) and the lattice translation have Lieb-Schultz-Mattis (LSM) anomaly. We show that $\gamma(\pi)=\ln 2$ (for even system size).
  2. 2.
    A translation-invariant spin-1/2 chain with an anomalous $\mathrm{O}(2)\times\mathds{Z}_{2}$ symmetry group, where the $\mathrm{U}(1)$ symmetry is non-on-site. We show that
[…]
While the ’t Hooft anomalies in these two systems appear quite different, they share a common feature: a $\pi$ flux of the U(1) symmetry carries a projective representation protected by the other symmetries. This type of ’t Hooft anomaly is often referred to as the “type-III” anomaly in literature de Wild Propitius (1995). The type-III structure of the anomaly turns out to be crucial for the quantization of $\gamma(\pi)$.
We compute $G(\theta)$ in two spin chain lattice models at finite temperature: 1. the spin-1/2 XX chain, which has the O(2) LSM anomaly, and 2. the Levin-Gu spin chain with the $\mathrm{O}(2)\times\mathds{Z}_{2}$ mixed anomaly. Both models can be solved exactly using Jordan-Wigner transformations, and the results for $\gamma(\pi)$ indeed agree with the expectations. We also match the full results for $G(\theta)$ with the CFT predictions at low energy, which requires the understanding of the continuum limit as the CFT with insertions of emanant symmetry defects. We also show explicitly in these examples that if the symmetry is reduced so the anomaly is no longer of type-III, $\gamma(\pi)$ is not quantized anymore and can change continuously depending on microscopic parameters, such as the filling factor.
More generally, we provide arguments for the quantization of $\gamma(\pi)$ based on the matrix product representation of the density matrix for both cases mentioned above, generalizing an argument in Zang et al. . As a by-product, we find that in both cases the U(1) symmetry can be reduced to the $\mathds{Z}_{2}$ subgroup without affecting the results, since the type-III anomaly structure is preserved.
II Charge fluctuations in a CFT
We first consider the case when the (1+1)d system is described by a CFT, at least at the energy scale of interest to us. More precisely, we assume that at low energy, the Hamiltonian $H$ can be approximated by
$H\approx\frac{2\pi v}{L}H_{\rm CFT}+E_{0}(L),$ (6)
where $L$ is the length of the system, $H_{\rm CFT}=L_{0}+\mkern 1.5mu\overline{\mkern-1.5muL\mkern-1.5mu}\mkern 1.5mu_{0}$ is the dimensionless CFT Hamiltonian on a unit circle $S^{1}$, and $v$ is the velocity. $E_{0}(L)$ is the ground state energy. We define a rescaled temperature $\tilde{\beta}=\frac{2\pi v\beta}{L}$.
In this case, we will show below that the $\gamma(\theta)$ can be related to vacuum degeneracy in the defect sector. To show this, recall that we need to evaluate
$G(\tilde{\beta},\theta)=\langle\,\mathrm{e}^{\mathrm{i}\theta Q}\rangle=\frac{\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\tilde{\beta}H_{\rm CFT}}}{\operatorname{Tr}\,\mathrm{e}^{-\tilde{\beta}H_{\rm CFT}}}=\frac{Z(\tilde{\beta},\theta)}{Z(\tilde{\beta},0)}.$ (7)
Modular invariance of the CFT implies
$Z(\tilde{\beta},0)=Z(1/\tilde{\beta},0).$ (8)
For the numerator, however, under modular transformation it becomes the partition function of the theory with a defect:
$Z(\tilde{\beta},\theta)=Z_{\theta}(1/\tilde{\beta}).$ (9)
Here we define
$Z_{\theta}(\tilde{\beta})=\operatorname{Tr}_{\theta}\,\mathrm{e}^{-\tilde{\beta}H_{\rm CFT}},$ (10)
where $\operatorname{Tr}_{\theta}$ means that we perform the trace in the Hilbert space of the CFT with a $\,\mathrm{e}^{\mathrm{i}\theta Q}$ defect. Therefore,
$G(\theta)=\frac{Z_{\theta}(1/\tilde{\beta})}{Z_{0}(1/\tilde{\beta})}.$ (11)
In the limit of large $L$ and thus small $\tilde{\beta}$, the modular transformed theory is then at low temperature, and the partition function can be approximated by keeping just the contribution from the ground state. We thus find
$G(\theta)\approx d_{\theta}\,\mathrm{e}^{-\frac{h_{\theta}}{\tilde{\beta}}}=d_{\theta}\,\mathrm{e}^{-\frac{h_{\theta}}{2\pi v\beta}L}.$ (12)
Here $d_{\theta}$ is the ground state degeneracy in the defect sector. $h_{\theta}$ is the scaling dimension of the $\mathrm{U}(1)$ defect. For a rational CFT, the presence of a $\mathrm{U}(1)$ symmetry implies that the CFT has $\mathrm{U}(1)_{k}\times\mathrm{U}(1)_{-{k}}$ Kac-Moody algebra, with $k$ being the level, and for $\theta\in[0,\pi]$, $h_{\theta}$ is given by the charged Cardy formula Dyer et al. (2018); Hosseini et al. (2020); Fan et al. (2023):
$h_{\theta}=k\theta^{2}.$ (13)
Therefore,
$\gamma(\theta)=\ln d_{\theta}.$ (14)
Now we recall that quite generally, anomaly in the $\mathrm{U}(1)$ symmetry implies that certain quantum number of the topological defect must change as $\theta$ continuously goes from 00 to $2\pi$. This is essentially equivalent to Laughlin’s flux-insertion argument. As a consequence, the ground state must become degenerate at some value of $\theta$ where levels cross. Thus, the generic behavior of $\gamma(\theta)$ is that it remains 0 except for a few isolated points where it jumps to a finite value.
We note that the argument can be easily generalized to other symmetries. In fact, Eq. (12) holds without any modification, except that $h_{\theta}$ should be replaced by the conformal weight of the corresponding defect operator. Essentially equivalent result was found in the context of topological disorder parameter in a (2+1)d gapped state in Chen et al. (2022).
A useful generalization is when the CFT is equipped with a topological defect. For this purpose, instead of a general CFT, we specialize to the example of a $c=1$ Luttinger liquid (or free compact boson CFT), arguably the simplest CFT with $\mathrm{U}(1)$ symmetry, which will be most relevant for our lattice model examples below. In general, such a CFT has two $\mathrm{U}(1)$ symmetries, generated by charges $Q_{m}$ and $Q_{w}$, respectively. The Hamiltonian has two parts: $H_{\rm CFT}=H_{0}+H_{\rm osc}$, where $H_{\rm osc}$ is the energy of the oscillator modes, which do not contribute to the global charge fluctuations. The “zero mode” energy $H_{0}$ takes the following form:
$H_{0}=\frac{1}{2}\left(\frac{Q_{m}^{2}}{R^{2}}+R^{2}Q_{w}^{2}\right).$ (15)
Here $R$ is the radius of the compact boson. In addition to the $\mathrm{U}(1)_{m}\times\mathrm{U}(1)_{w}$ symmetry, at a generic radius the CFT also has a charge-conjugation symmetry which acts as $Q_{m}\rightarrow-Q_{m},Q_{w}\rightarrow-Q_{m}$, so together the full symmetry is $[\mathrm{U}(1)_{m}\times\mathrm{U}(1)_{w}]\rtimes\mathds{Z}_{2}$.
Consider the CFT with a defect corresponding to $\,\mathrm{e}^{2\pi\mathrm{i}\eta_{m}Q_{m}}\,\mathrm{e}^{2\pi\mathrm{i}\eta_{w}Q_{w}}$, with $\eta_{m},\eta_{w}\in[0,1)$. The effect of the defect is to change the quantization conditions of the charges. The energy levels (the zero mode contributions) become
$H_{0}(\eta_{m},\eta_{w})=\frac{1}{2R^{2}}\left(Q_{m}+\eta_{w}\right)^{2}+\frac{R^{2}}{2}\left(Q_{w}+\eta_{m}\right)^{2}.$ (16)
Here let us consider the case $\eta_{m}=0$, and compute $\langle\,\mathrm{e}^{\mathrm{i}\theta Q_{m}}\rangle$. Details of the derivation can be found in Appendix B. Evaluating the partition functions, we find that
$\alpha(\theta)=\frac{R^{2}}{4\pi v\beta}[\theta]_{\pi}^{2},$ (17)
where we define
$[\theta]_{\pi}=\begin{cases}{\theta}&0\leq\theta\leq\pi\\ \theta-2\pi&\pi<\theta\leq 2\pi\end{cases}.$ (18)
As expected on general grounds, $\alpha(\theta)$ is “local” and insensitive to global boundary condition. Notice that $\alpha(\theta)$ exhibits a cusp at $\theta=\pi$, which appears to be a general feature of CFTs with a U(1) symmetry. A similar cusp was found in the coefficient of the leading term in the $\mathrm{U}(1)$ disorder parameter in a (1+1)d Luttinger liquid or a (2+1)d Fermi liquid Tan and Ryu (2020); Jiang et al. (2023); Cai and Cheng .
The full expression of $G(\theta)$ is found to be
$G(\theta)=\begin{cases}\,\mathrm{e}^{-\mathrm{i}\eta_{w}[\theta]_{\pi}}\,\mathrm{e}^{-\alpha(\theta)L}&\theta\neq\pi\\ 2\cos(\eta_{w}\pi)\,\mathrm{e}^{-\alpha(\pi)L}&\theta=\pi\end{cases}.$ (19)
The “universal” contribution $\gamma(\theta)$ is given by:
$\gamma(\theta)=\begin{cases}0&\theta\neq\pi\\ \ln\left\lvert 2\cos\eta_{w}\pi\right\rvert&\theta=\pi\end{cases}.$ (20)
Notice that in general $\left\lvert 2\cos\eta_{w}\pi\right\rvert$ is not an integer. For $\eta_{w}=\frac{1}{2}$, one finds that $\gamma(\pi)$ diverges, and $G(\pi)=0$.
Unlike $\alpha(\theta)$, the prefactor in $G(\theta)$ (i.e. $\gamma(\theta)$ and the phase factor $\omega(\theta)$) only depends on the defect parameter $\eta_{w}$, and has no dependence on other quantities that are sensitive to microscopic details, such as $v$ and $R^{2}$. In other words, $\gamma(\theta)$ and $\omega(\theta)$ appear to be robust against small changes to the theory.
However, an important caveat in this argument for the robustness of $\gamma$ and $\omega$ is the dependence of $\eta_{w}$ on the microscopic physics. It has been understood now that certain microscopic conditions, such as the filling factor, enter the low-energy theory as background topological defects required by anomaly matching Cheng et al. (2016); Metlitski and Thorngren (2018); Cho et al. (2017); Jian et al. (2018). As an example, consider a $c=1$ Luttinger liquid in a lattice system with $\mathrm{U}(1)$ filling factor $\nu$ (i.e. the average charge per unit cell is $\nu$). The low-energy physics of a system with periodic boundary condition of $N$ unit cells should be described by the Luttinger liquid with a defect $\eta_{w}=-N\nu$ Cheng and Seiberg (2023); Metlitski and Thorngren (2018); Else et al. (2021). Since the filling can be continuously tuned by applying a chemical potential, the defect parameter and thus $\gamma(\pi)$ can also be changed continuously.
It also happens in some cases that the filling is fixed by additional global symmetries (such as charge conjugation), in which case $\gamma(\pi)$ becomes quantized. Typically, in these cases there is an exact LSM-type ’t Hooft anomaly associated with the symmetries. For later references, we write down the expressions for $G(\theta)$ with the filling factor $\nu$:
$G(\theta)=\begin{cases}\,\mathrm{e}^{\mathrm{i}N\nu[\theta]_{\pi}}\,\mathrm{e}^{-\alpha(\theta)N}&\theta\neq\pi\\ 2\cos(\pi N\nu)\,\mathrm{e}^{-\alpha(\pi)N}&\theta=\pi\end{cases}.$ (21)
The results derived in this section assume that the spectrum is described by a CFT. For lattice models, CFT only describes the spectrum up to a certain energy scale (e.g. of the order of the bandwidth). One may wonder whether the results still hold when the temperature is comparable (or even higher) to the cut-off scale. In the next sections we turn to $G(\theta)$ in lattice models.
III Spin-1/2 chain with LSM anomaly
We now consider a familiar system, the spin-$1/2$ chain with $\mathrm{O}(2)=\mathrm{U}(1)\rtimes\mathds{Z}_{2}$ internal symmetry. The $\mathrm{U}(1)$ charge is given by:
$Q=\frac{1}{2}\sum_{n=1}^{N}\sigma^{z}_{n},$ (22)
and the $\mathds{Z}_{2}^{C}$ charge-conjugation symmetry generated by
$X=\prod_{n=1}^{N}\sigma^{x}_{n}$ (23)
Here $N$ is the number of sites. We will only consider spin chains with periodic boundary condition throughout this work.
Notice that the O(2) symmetry is on-site, so it is not anomalous on its own. However, because each spin-1/2 site carries a projective representation of the O(2) symmetry, the system exhibits a LSM anomaly between the O(2) and the lattice translation Cheng et al. (2016); Jian et al. (2018); Metlitski and Thorngren (2018); Else and Thorngren (2020).
Without the $\mathds{Z}_{2}$ symmetry, there is no ’t Hooft anomaly between the U(1) and lattice translation symmetry. However, if the total filling is fixed, e.g. $Q=0$ (corresponding to half filling in the hard-core boson basis), the system exhibits the “filling anomaly” for $\mathrm{U}(1)$ and lattice translation Oshikawa (2000); Cheng et al. (2016); Cho et al. (2017); Jian et al. (2018); Cheng and Seiberg (2023), forbidding the existence of a symmetric trivial state. However, unlike the O(2) LSM anomaly, the filling anomaly only appears in the subspace constrained to have a fixed filling. Therefore, in a grand canonical ensemble, where the total charge is allowed to fluctuate, the filling anomaly is absent.
A prototypical model in this system is the XX Hamiltonian:
$H=-\sum_{n=1}^{N}(\sigma_{n}^{x}\sigma_{n+1}^{x}+\sigma_{n}^{y}\sigma_{n+1}^{y}).$ (24)
For this model, $G(\theta)$ can be computed with Jordan-Wigner transformation. We will sketch the key ingredients for the extraction of $\gamma(\theta)$, and leave the full details to Appendix D. As it turns out, one needs to distinguish the case of $\mu=0$ (i.e. half filling) and $\mu\neq 0$.
$\mu=0$: half filling
First we consider the model with the full O(2) symmetry. Equivalently, the chemical potential is set to 00 in the partition function. We will express $G(\theta)$ in terms of the following quantities:
$Z_{ss^{\prime}}(\theta)=\prod_{j=\frac{s^{\prime}}{2}}^{N-1+\frac{s^{\prime}}{2}}\left(1+(-1)^{s}\,\mathrm{e}^{{\mathrm{i}\theta}}\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}}\right),$ (25)
where $s$ and $s^{\prime}$ take value in $0,1$. We have
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}=\\ \frac{1}{2}\left[Z_{00}(\theta)-Z_{10}(\theta)+Z_{01}(\theta)+Z_{11}(\theta)\right].$ (26)
To evaluate $Z_{ss^{\prime}}$, we use the Euler-MacLaurin formula to convert $\ln Z_{ss^{\prime}}$ into integrals. For example, for $\theta\neq\pi$:
$Z_{00}(\theta)\approx\,\mathrm{e}^{\frac{N}{2\pi}I_{1}\left(\theta,\beta\right)},$ (27)
where the integral $I_{1}(\theta,\beta)$ is defined in (104). Here $\approx$ means that there are $\mathcal{O}(N^{-1})$ corrections in the exponential.
[…]
$Z_{11}(0)=\prod_{j=\frac{1}{2}}^{N-\frac{1}{2}}(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}})$ $\approx 4\,\mathrm{e}^{\frac{N}{2\pi}I_{2}(\beta)}.$ (28)
The factor of $4$ is crucial for the calculation of $\gamma$.
Let us present the results. In the limit of large $N$, $G(\theta)$ takes the form given in Eq. (21) with $\nu=1/2$. In particular, for $N$ even we have
$\begin{split}\gamma(\theta)&=\begin{cases}0&\theta\neq\pi\\ \ln 2&\theta=\pi\end{cases}.\end{split}$ (29)
It is worth noting that these results are valid for any temperature $\beta\neq 0$. For $\beta=0$, we instead get $G(\theta)=\left(\cos\frac{\theta}{2}\right)^{N}$, which still gives $\gamma(\theta)=0$ for any $\theta\neq\pi$, but $G(\pi)=0$ so $\gamma(\pi)$ is not well-defined.
[…]
$H_{0}=Q_{m}^{2}+\frac{1}{4}\left(Q_{w}+\frac{1}{2}\right)^{2},$ (30)
and the ground states are two-fold degenerate: $Q_{m}=0$ and $Q_{w}=0,-1$. This degeneracy is guaranteed by the LSM anomaly, as shown directly in the lattice model in Appendix A.
More generally, we need to first consider the translation symmetry. Importantly, the lattice unit translation leads to an emanant $\mathds{Z}_{2}$ symmetry in the low-energy CFT. Namely, the lattice translation $T$ has the following representation in the low-energy theory:
$T=\,\mathrm{e}^{-\mathrm{i}\pi(Q_{m}+Q_{w})}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{L}P},$ (31)
where $P$ is the CFT momentum. The lattice system with $N$ sites in the continuum limit becomes the CFT with $\eta_{w}=\eta_{m}=-\frac{1}{2}N$, which then leads to Eq. (21) with $\nu=1/2$.
While the results can be understood within the low-energy theory, we emphasize that our derivation in fact applies to any temperature $\beta\neq 0$, even at high temperature when the system is presumably not described by a CFT.
$\mu\neq 0$
It is instructive to consider turning on a nonzero chemical potential $\mu$ (or a nonzero Zeeman field in the spin language), which results in a different ground state filling factor (or magnetization plateau) in the XX model:
$\nu=\frac{1}{\pi}\arccos\left(-\frac{\mu}{4}\right).$ (32)
Note that we choose the convention that $\nu=1/2$ for $\mu=0$.
[…]
$\left\lvert G(\theta)\right\rvert\approx\,\mathrm{e}^{-\alpha(\theta)N},$ (33)
so $\gamma(\theta)=0$. The phase factor of $G(\theta)$ now depends on $\beta$ and $\nu$ in a complicated way. Only when $\beta$ is large (i.e. low temperature), the phase factor approaches the CFT form $\,\mathrm{e}^{\mathrm{i}N\nu[\theta]_{\pi}}$.
For $\theta=\pi$ we are able to determine the phase factor as well:
$G(\pi)\approx 2\cos(\nu N\pi)\,\mathrm{e}^{-\alpha(\pi)N}.$ (34)
Notably, these results hold for any finite temperature, as long as $\beta>0$.
From the perspective of low-energy theory, the emanant symmetry from translation becomes $\,\mathrm{e}^{-2\pi\mathrm{i}\nu Q_{w}}$. A lattice system of $N$ sites flows to a U(1) free boson CFT with a defect $\eta_{w}=-N\nu$. Eq. (34) then immediately follows from Eq. (20).
This example suggests that $\gamma(\pi)$ is not quantized in general in a lattice system, as it varies smoothly with the chemical potential (or filling factor). However, in the presence of additional symmetry, such as charge conjugation symmetry that fixes the filling, then $\gamma(\pi)$ takes quantized value $\ln 2$ (for even $N$). A spin-1/2 chain with O(2) symmetry has the LSM anomaly, which is a true ’t Hooft anomaly that holds for the entire Hilbert space. For this reason we expect that the quantization of $\gamma(\pi)$ holds more generally. We show in the next subsection that this is indeed the case, and $\gamma(\pi)=\ln 2$ for even $N$ is a direct consequence of the LSM anomaly.
III.1 Quantization of $\gamma(\pi)$
We now show that $\gamma(\pi)=\ln 2$ holds for any thermal state of a local Hamiltonian (i.e. with short-range interactions), as long as the $\mathrm{O}(2)=\mathrm{U}(1)\rtimes\mathds{Z}_{2}^{C}$ symmetry is preserved. In fact, as will become clear all we need is the $\mathds{Z}_{2}\times\mathds{Z}_{2}$ subgroup, generated by $U_{Z}=\prod_{n}\sigma^{z}_{n}$ and $U_{X}=\prod_{n}\sigma^{x}_{n}$. We define
$G_{Z}=\operatorname{Tr}U_{Z}\rho.$ (35)
Note that $G(\pi)=\mathrm{i}^{N}G_{Z}$.
To proceed we follow the argument presented in Zang et al. . First we represent a translation-invariant density matrix as a matrix product operator (MPO) Verstraete et al. (2004); Perez-Garcia et al. (2007); Cirac et al. (2017, 2021):
$\rho$ =$=\sum_{\{s\},\{s^{\prime}\}}\operatorname{Tr}\left[\cdots\hat{M}^{s_{j}s^{\prime}_{j}}\hat{M}^{s_{j+1}s^{\prime}_{j+1}}\cdots\right]$ (36)
×$\qquad\qquad\times\lvert\dots s_{j}s_{j+1},\dots\rangle\langle\dots s^{\prime}_{j}s^{\prime}_{j+1}\dots\rvert$
[…]
It is known that such an approximation is always possible for thermal states of local Hamiltonians Hastings (2006); Molnar et al. (2015); Kuwahara et al. (2021); Huang (2021). Furthermore, since there is no long-range order at finite temperature in 1D, we expect that the MPO is injective Perez-Garcia et al. (2007); Cirac et al. (2021), i.e. the corresponding transfer matrix has a non-degenerate leading eigenvalue.
Using the MPO representation, we write
[…]
$=:\operatorname{Tr}\big{[}\big{(}{M}_{Z}\big{)}^{N}\big{]}.$
Here we denote $Z\equiv\sigma^{z}$, and define the “symmetry-twisted” transfer matrix ${M}_{Z}$ as
[…]
Crucially, the density matrix also commutes with the $X$ symmetry: $U_{X}\rho U_{X}=\rho$. Via the fundamental theorem of matrix product vectors, there must exist an invertible matrix $W_{X}$ such that
[…]
Applying this virtual symmetry to ${M}_{Z}$, it follows from the relation $XZX=-Z$ that
$W_{X}{M}_{Z}W^{-1}_{X}=-{M}_{Z}.$ (40)
In other words, $W_{X}$ and ${M}_{Z}$ anticommute. Therefore, all eigenvalues of ${M}_{Z}$ must come in pairs ±$\pm\lambda$, including those with the largest magnitude. In principle, it is possible that $\left\lvert\lambda\right\rvert$ has degeneracy more than 2. However, given that the leading eigenvalue of $M$ is non-degenerate, we expect that $M_{Z}$ has no more degeneracy in the leading $|\lambda|$ than what is required by the symmetry condition in Eq. (40). That is, the leading eigenvalues of ${M}_{Z}$ should be ±$\pm\lambda_{\rm max}$  [^1] It is possible that $\lambda_{\rm max}=0$, in which case $\gamma$ becomes meaningless. This is what happens at $\beta=0$. With this assumption,
$G_{Z}\approx\left[1+(-1)^{N}\right]\lambda_{\rm max}^{N}=\begin{cases}2\lambda_{\rm max}^{N}&N\text{ is even}\\ 0&N\text{ is odd}\end{cases}.$ (41)
From this result we immediately see that $\gamma(\pi)=\ln 2$.
Notice that the result applies to any system with $\mathds{Z}_{2}\times\mathds{Z}_{2}$ LSM anomaly Cheng et al. (2016); Po et al. (2017); Ogata and Tasaki (2019); Ogata et al. (2021), since the only symmetry property used here is $X_{n}Z_{n}=-Z_{n}X_{n}$. In fact, the only assumptions needed on $\rho$ are translation invariance and the $X$ symmetry.
Naturally one may wonder whether a similar argument can explain the $(-1)^{N/2}$ factor in $G(\pi)$ of the XX spin chain. We now make a further assumption that ${M}$ is hermitian (within a certain gauge), and $Z$ is real (i.e. has real matrix elements). In this case, we can easily show that ${M}_{Z}$ is also hermitian:
$\left[M_{Z}\right]_{ba}$ $=\sum_{ss^{\prime}}Z_{s^{\prime}s}{M}^{ss^{\prime}}_{ba}$ (42)
$=\sum_{ss^{\prime}}Z_{s^{\prime}s}\left({M}^{ss^{\prime}}_{ab}\right)^{*}$ (43)
$=\left(\sum_{ss^{\prime}}Z_{s^{\prime}s}{M}^{ss^{\prime}}_{ab}\right)^{*}$ (44)
$=\left[M_{Z}\right]_{ab}^{*}.$ (45)
As a result, the eigenvalues of ${M}_{Z}$ come in pairs ±$\pm|\lambda|$, and for even $N$ we have
$G_{Z}\approx 2\left\lvert\lambda_{\rm max}\right\rvert^{N}.$ (46)
While generally a transfer matrix may not be hermitian, it is known that e.g. the transfer matrix of the XXZ model
$H_{\rm XXZ}=\sum_{n=1}^{N}(\sigma_{n}^{x}\sigma_{n+1}^{x}+\sigma_{n}^{y}\sigma_{n+1}^{y}+\Delta\sigma^{z}_{n}\sigma^{z}_{n+1})$ (47)
can be made hermitian for $\Delta\geq 0$ Tang et al. (2020). Thus Eq. (46) applies to $H_{\rm XXZ}$ as well.
We note that the translation invariance is crucial for this argument, otherwise Eq. (36) would not hold. In fact, if the translation symmetry is broken (e.g. the unit cell is doubled) but the O(2) symmetry is preserved, one can easily construct examples with $\gamma(\theta)=0$ for all $\theta$.
IV Spin chain with mixed $\mathrm{O}(2)\times\mathds{Z}_{2}$ anomaly
In this section, we consider a spin chain with anomalous $\mathrm{O}(2)\times\mathds{Z}_{2}^{X}$ symmetry. Unlike the previous example, here the $\mathrm{U}(1)$ symmetry is non-on-site, and importantly there is a type-III ’t Hooft anomaly between $\mathds{Z}_{2}^{X}$ and O(2).
More specifically, the system is made of qubits, with the $\mathrm{U}(1)$ charge defined as
$Q=\frac{1}{4}\sum_{n}(1-\sigma_{n}^{z}\sigma_{n+1}^{z}).$ (48)
In the $\sigma^{z}$ eigenbasis, $2Q$ counts the number of domain walls, so with periodic boundary conditions, $Q$ takes integer values. Interestingly, $Q$ is manifestly non-on-site, but still anomaly-free. The constant term (equal to $N/4$) is included in $Q$ to ensure that $Q$ takes integer values for any system size $N$.
The charge-conjugation symmetry $\mathds{Z}_{2}^{C}$ in the O(2) group is generated by
$X_{\rm even}=\prod_{i=1}^{N/2}\sigma^{x}_{2i}.$ (49)
One can readily see that
$X_{\rm even}QX_{\rm even}^{-1}=\frac{1}{4}\sum_{n}(1+\sigma_{n}^{z}\sigma_{n+1}^{z})=-Q+\frac{N}{2}.$ (50)
Thus to preserve the O(2) symmetry the filling factor $\nu=Q/N$ must be $1/4$. We can similarly define $X_{\rm odd}$, which has the same action on $Q$.
Lastly, the on-site $\mathds{Z}_{2}^{X}$ symmetry is generated by
$X=\prod_{n}\sigma_{n}^{x}=X_{\rm even}X_{\rm odd}.$ (51)
Now we discuss the ’t Hooft anomaly of the internal $\mathds{Z}_{2}^{X}\times\mathrm{O}(2)$ symmetry. As alluded to above, there is a mixed anomaly between $\mathds{Z}_{2}^{X}$ and O(2). In fact, the same kind of anomaly is already present when restricted to $\mathds{Z}_{2}\subset\mathrm{U}(1)$. The $\mathds{Z}_{2}\times\mathds{Z}_{2}^{C}\times\mathds{Z}_{2}^{X}$ symmetry has an anomaly associated with the so-called “type-III” 3-cocycle de Wild Propitius (1995). Physically, it is characterized by the $\mathds{Z}_{2}$ defect transforming projectively under $\mathds{Z}_{2}^{\cal C}\times\mathds{Z}_{2}^{X}$.
Moreover, even the $\mathds{Z}_{2}^{C}$ symmetry is ignored, the remaining $\mathrm{U}(1)\times\mathds{Z}_{2}^{X}$ symmetry is still anomalous. The system can be viewed as a lattice model for the edge of a bosonic SPT state protected by the $\mathrm{U}(1)\times\mathds{Z}_{2}$ symmetry Chen et al. (2013); Metlitski (2019); Son and Alicea (2019); Horinouchi (2020). Turning on a non-zero chemical potential breaks $\mathds{Z}_{2}^{C}$ but preserves $\mathds{Z}_{2}^{X}$.
Below we will study the following Levin-Gu Hamiltonian preserving $\mathds{Z}_{2}^{X}\times\mathrm{O}(2)$:
$H_{\rm LG}=-\sum_{n=1}^{N}(\sigma_{n}^{x}-\sigma_{n-1}^{z}\sigma_{n}^{x}\sigma_{n+1}^{z}),$ (52)
which was first considered in Levin and Gu (2012) as an edge model of the nontrivial (2+1)d $\mathds{Z}_{2}$ SPT state. The model is exactly solvable, and can be mapped to a gauged XX spin chain as follows. Introduce a dual representation of the system Levin and Gu (2012):
[…]
where $\tau$ can be understood as domain wall variables and $\mu$ represent $\mathds{Z}_{2}$ gauge fields. They are subject to the Gauss’s law constraint:
$\tau_{n}^{z}=\mu_{n-1,n}^{x}\mu_{n,n+1}^{x}.$ (54)
In this representation, $Q$ can be written as
$Q$ $=\frac{1}{4}\sum_{n}(1-\tau_{n}^{z}).$ (55)
And similarly $X=\prod_{n}\mu_{n,n+1}^{x}$. The LG Hamiltonian becomes
$H_{\rm LG}=-\sum_{n=1}^{N}(\tau_{n}^{x}\tau_{n+1}^{x}+\tau_{n}^{y}\tau_{n+1}^{y})\mu_{n,n+1}^{z},$ (56)
namely a XX spin chain coupled to a $\mathds{Z}_{2}$ gauge field. $Q$ is the total spin z component and $X$ is the Wilson loop for the gauge field.
The partition function can be exactly evaluated by mapping to free fermions. We find that for any $\beta\geq 0$, $G(\theta)$ is given precisely by Eq. (21) with $\nu=1/4$. In particular, $\gamma$ is given by:
[…]
Details of the calculations can be found in Appendix C.
[…]
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}$ $=\operatorname{Tr}X_{\rm even}^{2}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}$ (58)
$=(-1)^{N/2}\operatorname{Tr}\,\mathrm{e}^{-\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}$ (59)
$=-\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}.$ (60)
Thus, we conclude that $G(\pi)=0$. Essentially, the system forms a nontrivial projective representation of $\mathrm{O}(2)$.
To understand the physical meaning of $\gamma$ and $\omega$, let us consider two limits. First, at low temperature $\beta\rightarrow\infty$, the system is described by a $c=1$ Luttinger liquid with $R=\frac{1}{\sqrt{2}}$ Cheng and Seiberg (2023). The $\mathrm{U}(1)$ charge $Q$ is identified with $Q_{m}$ in the low energy theory, and $X$ becomes $\,\mathrm{e}^{\mathrm{i}\pi Q_{w}}$. In addition, the lattice unit translation leads to an emanant $\mathds{Z}_{4}$ symmetry $\,\mathrm{e}^{-\frac{\mathrm{i}\pi}{2}Q_{w}}$ Cheng and Seiberg (2023). Namely, the lattice translation $T$ has the following representation in the low-energy theory:
$T=\,\mathrm{e}^{-\frac{\mathrm{i}\pi}{2}Q_{w}}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{L}P},$ (61)
where $P$ is the CFT momentum. As a result of the nontrivial emanant symmetry, a chain of size $N$ should flow to a CFT with a $\eta_{w}=-\frac{N}{4}$ defect. Therefore in this limit $G(\theta)$ takes the form of Eq. (21) with $\nu=1/4$.
[…]
We note in passing that adding a chemical potential (thus reducing O(2) to $\mathrm{U}(1)$) changes the ground state filling factor, and thus similar to the case of the XX model, the value of $\gamma(\pi)$ also changes continuously with the chemical potential.
The other limit is $\beta\rightarrow 0$, i.e. high temperature. In this case, we can neglect the Hamiltonian and compute the trace of the operator $\,\mathrm{e}^{\mathrm{i}\theta Q}$. This computation will be discussed in the next subsection.
In both limits, one finds that $\gamma(\pi)$ takes the quantized values given in as long as the $\mathds{Z}_{2}^{X}\times\mathds{Z}_{2}^{C}$ symmetry is preserved, which suggests that the quantization is independent of the details of the Hamiltonian. This will be established in Sec. IV.2.
IV.1 High temperature limit
It is instructive to consider the infinite-temperature state $\rho=\frac{1}{2^{N}}\mathds{1}$ first. Then
${G(\theta)}$ $=2^{-N}{\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}}$ (62)
$=2^{-N}\,\mathrm{e}^{\frac{\mathrm{i}\theta}{4}N}\operatorname{Tr}\,\mathrm{e}^{-\mathrm{i}\frac{\theta}{4}\sum_{n}\sigma_{n}^{z}\sigma_{n+1}^{z}}.$ (63)
Note that the trace is precisely the 1d Ising Hamiltonian with imaginary coupling. Generally we can write
$\operatorname{Tr}\exp\left(K\sum_{n=1}^{N}\sigma_{n}^{z}\sigma_{n+1}^{z}\right)=z_{+}^{N}+z_{-}^{N},$ (64)
where $z_{\pm}=\,\mathrm{e}^{K}\pm\,\mathrm{e}^{-K}$ are eigenvalues of the transfer matrix. In our case $K=-\frac{1}{4}\mathrm{i}\theta$, so we have
$2^{-N}\operatorname{Tr}\,\mathrm{e}^{-\frac{1}{4}\mathrm{i}\theta\sum_{n}\sigma_{n}^{z}\sigma_{n+1}^{z}}=\cos^{N}\frac{\theta}{4}+(-\mathrm{i})^{N}\sin^{N}\frac{\theta}{4}.$ (65)
First, let us look at the special case $\theta=\pi$. We have
$G(\pi)$ $=\,\mathrm{e}^{\frac{1}{4}\mathrm{i}\pi N}[1+(-\mathrm{i})^{N}]2^{-N/2}$ (66)
$=2\cos\left(\frac{\pi N}{4}\right)2^{-N/2}.$ (67)
For general $\theta\in(0,2\pi)$,
$\ln\left\lvert G(\theta)\right\rvert=\ln\left\lvert\cos^{N}\frac{\theta}{4}+(-\mathrm{i})^{N}\sin^{N}\frac{\theta}{4}\right\rvert.$ (68)
If $\theta<\pi$, then $\cos\frac{\theta}{4}>\sin\frac{\theta}{4}$, $\ln\left\lvert G(\theta)\right\rvert\approx N\ln\cos\frac{\theta}{4}$ for $N$ large. If $\pi<\theta<2\pi$, then $\sin\frac{\theta}{4}>\cos\frac{\theta}{4}$, $\ln\left\lvert G(\theta)\right\rvert\approx N\ln\sin\frac{\theta}{4}$ for $N$ large. In both cases, we have $\gamma=0$ as claimed.
We now consider a slightly deformed state, adding a chemical potential: $\rho\propto\,\mathrm{e}^{\lambda Q}$. Notice that $\lambda\neq 0$ breaks the $\mathds{Z}_{2}^{C}$ symmetry, but the $\mathds{Z}_{2}^{X}$ symmetry is preserved. It is straightforward to generalize the calculations above, and here we will just present the $\theta=\pi$ result:
$G(\pi)=2^{-\frac{N}{2}}\left[(1-\mathrm{i}\tanh\lambda)^{N}+(-\mathrm{i})^{N}(1+\mathrm{i}\tanh\lambda)^{N}\right],$ (69)
from which we can extract
$\gamma(\pi)=\ln\left\lvert 2\cos N\left(\varphi-\frac{\pi}{4}\right)\right\rvert,$ (70)
where $\varphi(\lambda)=\arctan(\tanh\lambda)$. One can also easily show that $\gamma(\theta)=0$ for $\theta\neq\pi$. Notice that in this state $\rho$, the U(1) charge density is given by $\frac{1}{4}(1-\tanh\lambda)$. This simple example shows that $\gamma(\pi)$ changes with the filling and is not quantized with just $\mathrm{U}(1)\times\mathds{Z}_{2}^{X}$.
IV.2 Quantization of $\gamma(\pi)$
In this section, we present arguments for the quantization of $\gamma(\pi)$ in this system. Before going to the details, it is important to clarify the role of translation invariance. According to Eq. (50), with full translation symmetry, the filling factor of the system is fixed at $1/4$, which is already a strong hint that $\gamma(\pi)$ should be universal. Even with a doubled unit cell, the filling factor is $1/2$ and one would expect that $\gamma(\pi)$ is quantized to $\ln 2$ just like the O(2) LSM case discussed in Sec. III.1 (even though the U(1) charge is not on-site, an important difference). However, we will find that the quantization $\gamma(\pi)=\ln 2$ does not really rely on LSM-type anomaly and is instead enforced by the anomalous internal symmetry. More specifically, we will show that the quantization holds with a four-site unit cell for which the filling factor is an integer. On the other hand, if the state has the full translation symmetry, $\gamma(\pi)$ exhibits interesting dependence on $N$ mod 4, which we will provide general arguments for.
First, because the symmetry operator $\,\mathrm{e}^{\mathrm{i}\pi Q}$ is non-on-site, it is represented as a MPO with the following tensor with bond dimension $D=2$ (up to an overall phase):
{tikzpicture}$\begin{tikzpicture}=\delta_{\alpha s}\mathrm{i}^{\alpha+\beta}(-1)^{\alpha\beta}.$ (71)
Here $\alpha,\beta=0,1$ are the bond indices.
It is straightforward to check that the tensor satisfies the following two conditions:
[…]
The first condition guarantees that $\,\mathrm{e}^{\mathrm{i}\pi Q}$ commutes with $X$. The second condition can be understood as a kind of “gauge symmetry” of the $U$ tensor.
$G(\pi)$ can be written as
$G(\pi)=\operatorname{Tr}{M}_{U}^{N},$ (74)
where the tensor ${M}_{U}$ is defined by the following diagram:
${M}_{U}:=\begin{tikzpicture}.$ (75)
Now we consider the $C$ symmetry. To this end, it is convenient to group two neighboring sites into a doubled unit cell. By our assumption, the density matrix commutes with $C$, and we must have
[…]
Here $V_{X}$ is an invertible matrix.
We now prove the following key relation:
$(Z\otimes V_{X})M_{U}^{2}=-M_{U}^{2}(Z\otimes V_{X}).$ (77)
It can be established by the following steps:
$=\mathrm{i}\begin{tikzpicture}$ (78)
$=\mathrm{i}\begin{tikzpicture}$ (79)
[…]
Here in the first step we apply the gauge symmetry condition to both $U$ tensors. From the first to the second line, we use the $C$ symmetry of the $M^{2}$ tensor given in Eq. (76). For the next step, we apply the $\mathds{Z}_{2}^{X}$ symmetry of the $U$ tensor on the right, and eliminate the remaining $X$ on the physical indices.
Therefore, all eigenvalues of $M_{U}^{2}$ must come in pairs ±$\pm\lambda$, including the largest one. Thus we have
$G(\pi)$ $\approx\left[1+(-1)^{N/2}\right]\lambda_{\rm max}^{N}$ (81)
[…]
From this result we immediately see that $\gamma(\pi)=\ln 2$.
Now if the largest eigenvalues of $M_{U}^{2}$ are ±$\pm\lambda_{\rm max}$, then for $M_{U}$ they must be $\lambda_{\rm max}^{\prime}$ and ±$\pm\mathrm{i}\lambda_{\rm max}^{\prime}$, where the $\pm$ sign can not be directly fixed by this argument. However, this sign ambiguity does not affect $\left\lvert G(\pi)\right\rvert$ and one finds that
$\gamma(\pi)=\ln 2\left\lvert\cos\frac{N\pi}{4}\right\rvert.$ (83)
In particular, $\gamma(\pi)=\ln\sqrt{2}$ for odd $N$.
[…]
Let us now show that the same holds assuming only $T^{4}$, which is beyond the LSM case. Denote by $S$ the transfer operator with four-site unit cells. Diagrammatically, $S$ can be represented as
$S=\begin{tikzpicture}.$ (84)
Here $M^{\prime}$ is the tensor of the MPO representation of $\rho$ with four sites in a unit cell. In the fully translation invariant case, we have $S=M_{U}^{4}$. Since $\rho$ is invariant under $X_{\rm even}$ ($X_{\rm odd}$), the action $X_{\rm even}$ ($X_{\rm odd}$) can be pushed to the virtual space, which will be denoted by $V_{\rm even}$ ($V_{\rm odd}$). Importantly, $V_{\rm even}$ and $V_{\rm odd}$ must commute, otherwise the transfer matrix obtained by contracting the physical indices of $M^{\prime}$ would have degenerate spectrum, contradicting the short-range nature of $\rho$.
Following steps very similar to those in Eq. (80), one can prove
$\begin{split}(Z\otimes V_{\rm even})S=S(Z\otimes V_{\rm even}),\\ (Y\otimes V_{\rm odd})S=S(Y\otimes V_{\rm odd}).\end{split}$ (85)
Basically, $S$ is invariant under the virtual $\mathds{Z}_{2}\times\mathds{Z}_{2}$ symmetry generated by $Z\otimes V_{\rm even}$ and $Y\otimes V_{\rm odd}$. Because $V_{\rm even}$ and $V_{\rm odd}$ commute, the virtual states of $S$ form a projective representation of $\mathds{Z}_{2}\times\mathds{Z}_{2}$ symmetry, and thus the eigenvalues of $S$ are at least two-fold degenerate. Again, generically we expect there is no further degeneracy in the spectrum of $S$, and with this assumption we obtain $\gamma(\pi)=\ln 2$.
Heuristically, the result follows from the fact that the unitary $\,\mathrm{e}^{\mathrm{i}\pi Q}$ is a $\mathds{Z}_{2}\times\mathds{Z}_{2}$ SPT entangler.
V Conclusions and Discussions
In this work we study the constant correction $\gamma(\theta)$ in the thermal expectation value $G(\theta)$ of $\,\mathrm{e}^{\mathrm{i}\theta Q}$, for a (1+1)d periodic lattice system with U(1) symmetry. We show that $\gamma(\pi)$ becomes quantized in two systems where the U(1) symmetry has a type-III mixed anomaly with other global symmetry of the system. The value of $\gamma(\pi)$ is closely linked to the symmetry-protected degeneracy of U(1) symmetry defect. Without such type of anomaly, $\gamma(\pi)$ can depend on microscopic parameters, such as the filling factor. We also provide field-theoretical understandings of these results when the system can be described by a CFT.
An important question left for future work is to more systematically study the relation between the quantization of $\gamma$ and type-III ’t Hooft anomaly, beyond the specific example considered in this work. From the CFT perspective, the universal contribution $\gamma$ for a given symmetry operator $g$ is given by $\ln d_{g}$, where $d_{g}$ is the degeneracy of the $g$ defect. If $g$ has a type-III mixed anomaly, it means that the $g$ defect transforms projectively under the remaining symmetry $Z_{g}$ (the centralizer of $g$). The projective class is determined by the anomaly 3-cocycle, as explained in Appendix A. We conjecture that generically $d_{g}$ is the minimal dimension of the irreducible representation in the same projective class.
Following this line of thought, one expects that similar results should hold for fermionic systems. For example, in a (1+1)d fermionic system with $\mathds{Z}_{2}\times\mathds{Z}_{2}^{F}$ symmetry ($\mathds{Z}_{2}^{F}$ stands for the fermion parity conservation), the ’t Hooft anomaly is classified by $\mathds{Z}_{8}$. The generator of the $\mathds{Z}_{8}$ is characterized by the $\mathds{Z}_{2}$ symmetry defect carrying a Majorana zero mode. Thus we expect that if one measures the expectation value of the total $\mathds{Z}_{2}$ charge in a thermal state, $\gamma$ should take a universal value $\gamma=\ln\sqrt{2}$. Another system with a mathematically similar anomaly is a translation-invariant chain of Majorana modes Cheng (2019), where the lattice translation has a mixed anomaly with fermion parity. In this case, we expect that the thermal expectation value of the translation operator contains a universal correction $\ln\sqrt{2}$.
An obvious direction for future works is to generalize the results to higher dimensions. It is not difficult to see that the MPO argument for $\gamma(\pi)=\ln 2$ can be generalized to two dimensions, assuming a PEPO representation of the thermal density matrix. It will be worth investigating other classes of systems, such as Fermi liquid or quantum critical points, or systems with other types of ’t Hooft anomalies.
VI Acknowledgements
We thank Yingfei Gu, Shenghan Jiang and Pengfei Zhang for helpful conversations and communications on their unpublished works. M.C. is particularly grateful for Hong-Hao Tu for collaboration on a previous related project and many enlightening discussions on tensor networks. M.C. would like to acknowledge NSF for support under award number DMR-1846109.
Note added: During the finalization of the manuscript, we became aware of closely related works Zang et al. and Wang et al. , which study the universal feature in $\alpha(\theta)$. In particular, Zang et al. showed that the O(2) LSM anomaly leads to the cusp at $\theta=\pi$ in $\alpha(\theta)$.
Appendix A Typology of ’t Hooft anomaly in (1+1)d
In (1+1)d bosonic/spin systems, ’t Hooft anomaly for a global symmetry $G$ is classified by $\mathcal{H}^{3}(G,\mathrm{U}(1))$. Namely, each anomaly is uniquely associated with a group cohomology class $[\omega]\in\mathcal{H}^{3}(G,\mathrm{U}(1))$.
Anomalies can be partially characterized by symmetry transformation properties of a symmetry defect. More concretely, given $g\in G$, one considers the system with a $g$ defect, which can be viewed as a (0+1)d quantum-mechanical system with $Z_{g}$ symmetry, where $Z_{g}=\{h\in G|hg=gh\}$ is the centralizer of $g$. The $Z_{g}$ symmetry action may be projective, characterized by a 2-cocycle $\omega_{g}=i_{g}\omega$ in $\mathcal{H}^{2}(Z_{g},\mathrm{U}(1))$, where $i_{g}$ is the slant product. The explicit expression for $i_{g}\omega$ is given by
$(i_{g}\omega)(h,k)=\frac{\omega(g,h,k)\omega(h,k,g)}{\omega(h,g,k)}.$ (86)
If $\omega_{g}$ is nontrivial for some $g\in G$, the anomaly $\omega$ is said to be type-III. To give an example, consider $G=\mathds{Z}_{2}^{3}$. Label the group elements by $a\equiv(a_{1},a_{2},a_{3})$ where $a_{1},a_{2},a_{3}\in\{0,1\}$, and the group multiplication is defined as addition mod 2. The type-III cocycle is given by
$\omega(a,b,c)=(-1)^{a_{1}b_{2}c_{3}}.$ (87)
To see that it is indeed type-III, we compute $i_{(1,0,0)}$:
$(i_{(1,0,0)}\omega)(b,c)=(-1)^{b_{2}c_{3}}.$ (88)
This is the 2-cocycle for the projective representation of the $\mathds{Z}_{2}^{2}$ subgroup generated by $(0,1,0)$ and $(0,0,1)$.
Let us show that the two examples discussed in the main text have type-III anomaly.
We start with the XX spin chain. The Hamiltonian can be modified to have a $\theta=\pi$ defect at the link $N,1$:
$\begin{split}H(\pi)=-\sum_{n=1}^{N-1}(\sigma_{n}^{x}\sigma_{n+1}^{x}+\sigma_{n}^{y}\sigma_{n+1}^{y})+\sigma_{N}^{x}\sigma_{1}^{x}+\sigma_{N}^{y}\sigma_{1}^{y}.\end{split}$ (89)
It is clear that $X$ remains a symmetry of $H(\pi)$, but the lattice translation needs to be modified:
$T(\pi)=\,\mathrm{e}^{\mathrm{i}\frac{\pi}{2}\sigma^{z}_{1}}T=\mathrm{i}\sigma_{1}^{z}T.$ (90)
Therefore $T(\pi)X=-XT(\pi)$, implying that the system with a $\mathds{Z}_{2}\subset\mathrm{U}(1)$ defect transforms projectively under the $\mathds{Z}_{2}^{C}$ and lattice translation.
Next we consider the Levin-Gu model, introducing a $\theta=\pi$ defect Cheng and Seiberg (2023):
[…]
Apparently, $X=\prod_{n}\mu_{n,n+1}^{z}$ remains a symmetry for $H_{\rm LG}(\pi)$. However, the last term changes sign under $X_{\rm even}$. Therefore we need to redefine $X_{\rm even}$ as
$X^{\prime}_{\rm even}=\mu^{x}_{N,1}X_{\rm even},$ (92)
which then implies $X^{\prime}_{\rm even}X=-XX^{\prime}_{\rm even}$.
Appendix B $G(\theta)$ of the $c=1$ free boson CFT
Suppose the total charge $Q=t_{m}Q_{m}+t_{w}Q_{w}$, then
$G(\theta)=\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H_{0}(\eta_{m},\eta_{w})}=G_{m}(\theta_{m})G_{w}(\theta_{w}).$ (93)
Here we have defined $\theta_{m}=t_{m}\theta$ and $\theta_{w}=t_{w}\theta$. We have factorized the sum over $Q_{m}$ and $Q_{w}$, so in the following we only consider $G_{m}(\theta_{m})$. More explicitly:
$G_{m}(\theta_{m})=\frac{\sum_{Q_{m}\in\mathbb{Z}}\,\mathrm{e}^{\mathrm{i}\theta_{m}Q_{m}}\,\mathrm{e}^{-\frac{\beta}{2R^{2}}(Q_{m}+\eta_{w})^{2}}}{\sum_{Q_{m}\in\mathbb{Z}}\,\mathrm{e}^{-\frac{\beta}{2R^{2}}(Q_{m}+\eta_{w})^{2}}}=\frac{\vartheta\left(\frac{\theta_{m}}{2\pi}+\frac{\mathrm{i}\beta\eta_{w}}{2\pi R^{2}},\frac{\mathrm{i}\beta}{2\pi R^{2}}\right)}{\vartheta\left(\frac{\mathrm{i}\beta\eta_{w}}{2\pi R^{2}},\frac{\mathrm{i}\beta}{2\pi R^{2}}\right)}.$ (94)
The theta function is defined as
$\vartheta(z,\tau)=\sum_{n\in\mathbb{Z}}\,\mathrm{e}^{\mathrm{i}\pi\tau n^{2}+2\pi\mathrm{i}zn}.$ (95)
Using the S transformation
$\vartheta(z,\tau)=\frac{1}{\sqrt{-\mathrm{i}\tau}}\,\mathrm{e}^{-\mathrm{i}\pi\frac{z^{2}}{\tau}}\vartheta\left(\frac{z}{\tau},-\frac{1}{\tau}\right),$ (96)
we find
$\frac{\vartheta\left(\frac{\theta_{m}}{2\pi}+\frac{\mathrm{i}\beta\eta_{w}}{2\pi R^{2}},\frac{\mathrm{i}\beta}{2\pi R^{2}}\right)}{\vartheta\left(\frac{\mathrm{i}\beta\eta_{w}}{2\pi R^{2}},\frac{\mathrm{i}\beta}{2\pi R^{2}}\right)}$ $=\,\mathrm{e}^{-\mathrm{i}\theta_{m}\eta_{w}}\,\mathrm{e}^{-\frac{R^{2}\theta_{m}^{2}}{2\beta}}\frac{\sum_{n}\exp\left(-\frac{2\pi^{2}R^{2}}{\beta}(n^{2}-\frac{\theta_{m}}{\pi}n)+2\pi\mathrm{i}\eta_{w}n\right)}{\sum_{n}\exp\left(-\frac{2\pi^{2}R^{2}}{\beta}n^{2}+2\pi\mathrm{i}\eta_{w}n\right)}$ (97)
$\approx\,\mathrm{e}^{-\mathrm{i}\theta_{m}\eta_{w}}\,\mathrm{e}^{-\frac{R^{2}\theta_{m}^{2}}{2\beta}}{\sum_{n}\exp\left(-\frac{2\pi^{2}R^{2}}{\beta}(n^{2}-\frac{\theta_{m}}{\pi}n)+2\pi\mathrm{i}\eta_{w}n\right)}$ (98)
In the second line, we consider the $\beta\rightarrow 0$ limit and keep only the leading terms in the sum.
For $0\leq\theta_{m}<\pi$, we have
$\sum_{n}\exp\left(-\frac{2\pi^{2}R^{2}}{\beta}(n^{2}-\frac{\theta_{m}}{\pi}n)+2\pi\mathrm{i}\eta_{w}n\right)\approx 1,$ (99)
and for $\pi<\theta_{m}<2\pi$
$\sum_{n}\exp\left(-\frac{2\pi^{2}R^{2}}{\beta}(n^{2}-\frac{\theta_{m}}{\pi}n)+2\pi\mathrm{i}\eta_{w}n\right)\approx\,\mathrm{e}^{2\pi\mathrm{i}\eta_{w}}\exp\left(\frac{2\pi R^{2}(\theta_{m}-\pi)}{\tilde{\beta}}\right),$ (100)
and for $\theta_{m}=\pi$
$\sum_{n}\exp\left(-\frac{2\pi^{2}R^{2}}{\beta}(n^{2}-n)+2\pi\mathrm{i}\eta_{w}n\right)\approx 1+\,\mathrm{e}^{2\pi\mathrm{i}\eta_{w}}=2\,\mathrm{e}^{\mathrm{i}\pi\eta_{w}}\cos\pi\eta_{w}.$ (101)
Therefore, we find (for $\eta_{m}\neq 1/2$ mod $1$):
$\begin{split}G(\theta_{m})\approx\,\mathrm{e}^{-\mathrm{i}[\theta_{m}]_{\pi}\eta_{w}}\,\mathrm{e}^{-\frac{1}{2\beta}R^{2}[\theta_{m}]^{2}},\theta_{m}\neq\pi\\ G(\pi)\approx 2\cos(\eta_{w}\pi)\,\mathrm{e}^{-\frac{1}{2\beta}\pi^{2}R^{2}},\theta_{m}=\pi.\end{split}$ (102)
Appendix C $G(\theta)$ of the Levin-Gu model
First we define a few quantities that appear frequently in the calculations below. $Z_{ss^{\prime}}(\phi,\beta)$ is defined as
$Z_{ss^{\prime}}(\phi,\beta)=\prod_{j=\frac{s^{\prime}}{2}}^{N-1+\frac{s^{\prime}}{2}}\left(1+(-1)^{s}\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}}\right).$ (103)
Essentially, $Z_{ss^{\prime}}$ is the partition function of free fermions with periodic ($s^{\prime}=0$) or anti-periodic ($s^{\prime}=1$) boundary conditions, projected to the fermion parity $(-1)^{s}$ sector.
We will approximate the discrete product in $Z_{ss^{\prime}}$ by integrals. Define
$I_{1}(\phi,\beta)$ $=\int_{0}^{2\pi}\mathop{}\!\mathrm{d}x\,\ln\left(1+\,\mathrm{e}^{\mathrm{i}{\phi}}\,\mathrm{e}^{4\beta\cos x}\right),$ (104)
$I_{2}(\beta)$ $=\int_{0}^{2\pi}\ln\left\lvert 1-\,\mathrm{e}^{4\beta\cos x}\right\rvert\mathop{}\!\mathrm{d}x.$ (105)
A crucial property of $I_{1}(\phi,\beta)$ for our derivations is that for $-\pi<\phi<\pi$,
$\Im I_{1}(\phi,\beta)=\pi\phi.$ (106)
We show below in Appendix E that for $\phi\neq\pi$, we have
$\sum_{j=\frac{s^{\prime}}{2}}^{N-1+\frac{s^{\prime}}{2}}\ln\left(\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}+1\right)=\frac{N}{2\pi}\int_{0}^{2\pi}\mathop{}\!\mathrm{d}x\,\ln(\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{\alpha\cos x}+1)+\mathcal{O}(N^{-1}).$ (107)
As a result,
$Z_{0s^{\prime}}(\phi,\beta)\approx\exp\left[\frac{N}{2\pi}I_{1}(\phi,\beta)\right]=\,\mathrm{e}^{\frac{1}{2}\mathrm{i}[\phi]_{\pi}N}\exp\left[\frac{N}{2\pi}\Re I_{1}(\phi,\beta)\right].$ (108)
[…]
$\sum_{j=\frac{1}{2}}^{\frac{N}{4}-\frac{1}{2}}\ln(\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}-1)=\frac{N}{2\pi}\int_{0}^{\frac{\pi}{2}}\mathop{}\!\mathrm{d}x\,\ln(\,\mathrm{e}^{\alpha\cos x}-1)+\ln\sqrt{2}+\mathcal{O}(N^{-2}).$ (109)
We also need the following inequalities between the integrals. First note that
$\left\lvert 1+\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{4\beta\cos x}\right\rvert^{2}=1+\,\mathrm{e}^{8\beta\cos x}+2\cos\phi\,\mathrm{e}^{4\beta\cos x}\leq(1+\,\mathrm{e}^{4\beta\cos x})^{2},$ (110)
which implies that
$I_{1}(0,\beta)\geq\Re I_{1}\left(\phi,\beta\right).$ (111)
Similarly
$\,\mathrm{e}^{4\beta\cos x}+1>\left\lvert\,\mathrm{e}^{4\beta\cos x}-1\right\rvert\Longrightarrow I_{1}(0,\beta)>I_{2}(\beta).$ (112)
C.1 N even
For $N$ even, the Levin-Gu model can be mapped to a free fermion system with even number of fermions and periodic or anti-periodic boundary condition, we have
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}=$ $\frac{\,\mathrm{e}^{\frac{1}{2}\mathrm{i}\theta N}}{2}\left[Z_{00}(-\theta/2,\beta)+Z_{10}(-\theta/2,\beta)+Z_{01}(-\theta/2,\beta)+Z_{11}(-\theta/2,\beta)\right].$ (113)
First we consider $0<\theta<\pi$. As shown in Eq. (108),
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}$ $\approx\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]+\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2}+\pi,\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (114)
$\approx\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (115)
$=\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right].$ (116)
The second line follows from $\left\lvert 1+\,\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}}\,\mathrm{e}^{4\beta\cos x}\right\rvert>\left\lvert 1-\,\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}}\,\mathrm{e}^{4\beta\cos x}\right\rvert$ for any $0<\theta<\pi$. From this result we immediately see $\omega(\theta)=\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)$ for $0<\theta<\pi$.
For $\theta=\pi$, from (108) we have
$Z_{0s^{\prime}}(\pm\pi/2,\beta)=Z_{1s^{\prime}}(\mp\pi/2,\beta)=\exp\left[\frac{N}{2\pi}I_{1}\left(\pm\frac{\pi}{2},\beta\right)\right].$ (117)
Crucially, the integrals $I_{1}(\pm\frac{\pi}{2},\beta)$ satisfy $\Re I_{1}(\pm\frac{\pi}{2},\beta)=I(\beta)$, where $I(\beta)$ is given explicitly by
$I(\beta)=\int_{0}^{2\pi}\ln\sqrt{1+\,\mathrm{e}^{8\beta\cos x}}\mathop{}\!\mathrm{d}x.$ (118)
Therefore we have
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}\approx 2\cos\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}I(\beta)\right].$ (119)
For the partition function
$\operatorname{Tr}\,\mathrm{e}^{-\beta H}=$ $\frac{1}{2}\left[Z_{00}(0,\beta)+Z_{10}(0,\beta)+Z_{01}(0,\beta)+Z_{11}(0,\beta)\right],$ (120)
[…]
$Z_{10}(0,\beta)=\prod_{j=0}^{N-1}(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}})=0,$ (121)
and by Eq. (109)
$Z_{11}(0,\beta)$ $=\prod_{j=\frac{1}{2}}^{N-\frac{1}{2}}(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}})=\left[\prod_{j=\frac{1}{2}}^{\frac{N}{4}-\frac{1}{2}}(\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}}-1)\right]^{2}\left[\prod_{j=\frac{N}{4}+\frac{1}{2}}^{\frac{N}{2}-\frac{1}{2}}(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}})\right]^{2}$ (122)
$\approx 4\exp\left[\frac{N}{2\pi}I_{2}(\beta)\right].$ (123)
[…]
Putting everything together, we have found that for $0<\theta<\pi$,
$G(\theta)\approx\frac{\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right]}{\exp\left[\frac{N}{2\pi}I_{1}(0,\beta)\right]+2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}I_{2}(\beta)\right]}\approx\,\mathrm{e}^{\frac{1}{4}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right]},$ (124)
and for $\theta=\pi$,
$G(\pi)\approx\frac{2\cos\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}I(\beta)\right]}{\exp\left[\frac{N}{2\pi}I_{1}(0,\beta)\right]+2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}I_{2}(\beta)\right]}\approx 2\cos\left(\frac{\pi N}{4}\right)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-I(\beta)\right]}.$ (125)
C.2 N odd
When $N$ is odd, the Levin-Gu model corresponds to a free fermion system with odd number of fermions, and we can write
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}=\frac{\,\mathrm{e}^{\frac{1}{2}\mathrm{i}\theta N}}{2}\left[Z_{00}(-\theta/2,\beta)-Z_{10}(-\theta/2,\beta)+Z_{01}(-\theta/2,\beta)-Z_{11}(-\theta/2,\beta)\right].$ (126)
For $0<\theta<\pi$, we have
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}$ $\approx\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]-\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2}+\pi,\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (127)
$\approx\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (128)
$=\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right],$ (129)
while for $\theta=\pi$, we have
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}\approx$ $\left[\exp\left(\frac{\mathrm{i}\pi N}{4}\right)-\exp\left(\frac{3\mathrm{i}\pi N}{4}\right)\right]\exp\left(\frac{N}{2\pi}I(\beta)\right)$ (130)
$=$ $2(-1)^{\frac{N-1}{2}}\sin\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}I(\beta)\right].$ (131)
The large $N$ behavior of the partition function is still governed by $I_{1}(0,\beta)$ as $I_{1}(0,\beta)>I_{2}(\beta)$. Moreover, the two terms $Z_{10}(0,\beta)$ and $Z_{11}(0,\beta)$ have opposite sign for $N$ odd and they cancel each other. Write $N=4m+r$ where $r=1,3$, then $\sin\frac{\pi N}{4}=\sin(m\pi+\frac{r\pi}{4})=(-1)^{m}\sin\frac{r\pi}{4}=\frac{1}{\sqrt{2}}(-1)^{m}$. We conclude that for $0<\theta<\pi$,
$G(\theta)\approx\frac{\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right]}{\exp\left[\frac{N}{2\pi}I_{1}(0,\beta)\right]}=\,\mathrm{e}^{\frac{1}{4}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right]},$ (132)
and for $\theta=\pi$,
$G(\pi)\approx\frac{(-1)^{m+\frac{r-1}{2}}\sqrt{2}\exp\left[\frac{N}{2\pi}I(\beta)\right]}{\exp\left[\frac{N}{2\pi}I_{1}(0,\beta)\right]}=(-1)^{m+\frac{r-1}{2}}\sqrt{2}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-I(\beta)\right]}.$ (133)
Notice that $\cos\frac{\pi N}{4}=\cos(m\pi+\frac{\pi r}{4})=(-1)^{m}\cos\frac{\pi r}{4}=(-1)^{m}(-1)^{\frac{r-1}{2}}\frac{1}{\sqrt{2}}$. So the phase factor $(-1)^{m+\frac{r-1}{2}}$ is equal to $\operatorname{sgn}(\cos\frac{\pi N}{4})$ in the main text.
Appendix D $G(\theta)$ for XX spin chain
D.1 Half filling
We can write $\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}$ as
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}=\frac{1}{2}\left[Z_{00}(\theta,\beta)-Z_{10}(\theta,\beta)+Z_{01}(\theta,\beta)+Z_{11}(\theta,\beta)\right].$ (134)
For $0<\theta<\pi$, we have
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}\approx\exp\left[\frac{N}{2\pi}I_{1}\left(\theta,\beta\right)\right]=\exp\left(\frac{1}{2}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\theta,\beta\right)\right].$ (135)
For $\theta=\pi$,
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}=\frac{1}{2}\left[-Z_{00}(0,\beta)+Z_{10}(0,\beta)+Z_{01}(0,\beta)+Z_{11}(0,\beta)\right].$ (136)
We have proved that $\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}=0$ for $N$ odd, so we only need to consider $N$ even case. $Z_{00}(0,\beta)$ and $Z_{01}(0,\beta)$ can be approximated by Eq. (108), and they cancel each other. The other two terms are already considered in Appendix C, and we have $\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}=2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}I_{2}(\beta)\right]$. Note the prefactor $2$, which arises from $\mathcal{O}\left(\frac{1}{N}\right)$ correction to approximating the sum by an integral, is key to the topological correction $\gamma$ in XX spin chain.
The partition function $\operatorname{Tr}\,\mathrm{e}^{-\beta H}$ contains the same terms that appear in $\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}$ with different signs.
In conclusion, for $0<\theta<\pi$, we have
$G(\theta)\approx\,\mathrm{e}^{\frac{1}{2}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-\Re I_{1}(\theta,\beta)\right]}.$ (137)
While for $\theta=\pi$,
$G(\pi)$ $\approx 2\cos\left(\frac{N\pi}{2}\right)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-I_{2}(\beta)\right]}.$ (138)
D.2 Away from half filling
With a nonzero chemical potential $\mu$, the partition function of XX spin chain can be written as
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}=$ $\frac{1}{2}\left[\prod_{j=0}^{N-1}\left(1+\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}-4\beta\cos(\pi\nu)+\mathrm{i}\theta}\right)-\prod_{j=0}^{N-1}\left(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}-4\beta\cos(\pi\nu)+\mathrm{i}\theta}\right)\right]+$
$\frac{1}{2}\left[\prod_{j=\frac{1}{2}}^{N-\frac{1}{2}}\left(1+\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}-4\beta\cos(\pi\nu)+\mathrm{i}\theta}\right)+\prod_{j=\frac{1}{2}}^{N-\frac{1}{2}}\left(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}-4\beta\cos(\pi\nu)+\mathrm{i}\theta}\right)\right],$ (139)
where $\nu=\frac{1}{\pi}\arccos\left(-\frac{\mu}{4}\right)$ is the filling fraction.
We can approximate the products by integrals as
$\operatorname{Tr}\,\mathrm{e}^{-\beta H}\approx$ $\exp\left[\frac{N}{2\pi}I_{1}(0,\beta,\nu)\right],$ (140)
where
$I_{1}(\phi,\beta,\nu)$ $=\int_{0}^{2\pi}\ln\left(1+\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{4\beta\cos x-4\beta\cos(\pi\nu)}\right)\mathop{}\!\mathrm{d}x.$ (141)
And for $0<\theta<\pi$, we have
$\left\lvert\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}\right\rvert\approx\exp\left[\frac{N}{2\pi}\Re I_{1}(\theta,\beta,\nu)\right].$ (142)
The phase factor now depends on all the parameters $\theta,\beta$ and $\nu$. For $\theta=\pi$, define
$I_{2}(\beta,\nu)$ $=\int_{0}^{2\pi}\ln\left\lvert 1-\,\mathrm{e}^{4\beta\cos x-4\beta\cos(\pi\nu)}\right\rvert\mathop{}\!\mathrm{d}x.$ (143)
We show in Appendix E that the constant correction relating $I_{2}(\beta,\nu)$ and $\ln\prod\left(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}-4\beta\cos(\pi\nu)}\right)$ is now $2\ln\left[2\sin(\pi\delta)\right]$, where $\delta$ (or $1-\delta$) is the minimal distance between $j$ and the singular point $x=\frac{N\nu}{2}$. For $j$ taking integer values, we can take $\delta=\left[\frac{N\nu}{2}\right]$, which is the decimal part of $\frac{N\nu}{2}$. For $j$ taking half-integer values, we can take $\delta=\left\lvert\left[\frac{N\nu}{2}\right]-\frac{1}{2}\right\rvert$. Thus,
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}$ $\approx-2\sin^{2}\left(\left[\frac{N\nu}{2}\right]\pi\right)\exp\left[I_{2}(\beta,\nu)\right]+2\sin^{2}\left(\left[\frac{N\nu}{2}\right]\pi-\frac{1}{2}\pi\right)\exp\left[I_{2}(\beta,\nu)\right]$ (144)
$=2\cos\left(\nu N\pi\right)\exp\left[\frac{N}{2\pi}I_{2}(\beta,\nu)\right].$ (145)
In conclusion, for $0<\theta<\pi$, we have
$\left\lvert G(\theta)\right\rvert\approx\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta,\nu)-\Re I_{1}(\theta,\beta,\nu)\right]}.$ (146)
While for $\theta=\pi$,
$G(\pi)\approx 2\cos(\nu N\pi)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta,\nu)-I_{2}(\beta,\nu)\right]}.$ (147)
Appendix E Evaluation of the product
The Euler-MacLaurin formula for a continuously differentiable function $f(x)$ reads
$\sum_{i=a}^{b}f(i)=\int_{a}^{b}f(x)\mathop{}\!\mathrm{d}x+\frac{f(a)+f(b)}{2}+R_{m}.$ (148)
The remainder term $R_{m}$ is given by
$R_{m}=\sum_{k=1}^{m}\frac{B_{2k}}{(2k)!}[f^{(2k-1)}(b)-f^{(2k-1)}(a)]+\int_{a}^{b}\mathop{}\!\mathrm{d}x\,P_{2m+1}(x)f^{(2m+1)}(x),\>m\in\mathds{N}.$ (149)
Here $B_{k}$ are the Bernoulli numbers, and the periodized Bernoulli functions $P_{k}(x)$ is defined as $P_{k}(x)=B_{k}(x-[x])$, where $B_{k}(x)$ is the Bernoulli polynomial.
A particularly useful result is that when $|f^{\prime}(x)|$ is bounded on $[a,b]$:
$\left\lvert R_{0}\right\rvert=\left\lvert\int_{a}^{b}\mathop{}\!\mathrm{d}x\,P_{1}(x)f^{\prime}(x)\right\rvert\leq\int_{a}^{b}\mathop{}\!\mathrm{d}x\,\left\lvert P_{1}(x)f^{\prime}(x)\right\rvert\leq\int_{a}^{b}\mathop{}\!\mathrm{d}x\,\left\lvert f^{\prime}(x)\right\rvert\leq(b-a)\max_{x\in[a,b]}|f^{\prime}(x)|.$ (150)
Here we have used $\left\lvert P_{1}(x)\right\rvert=\left\lvert x-[x]-\frac{1}{2}\right\rvert\leq 1$.
We now use the Euler-MacLaurin formula to study the sum
$\sum_{j=0}^{N-1}\ln\left(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}+1\right).$ (151)
Applying Eq. (148) with $f(x)=\ln\left(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos\frac{2\pi x}{N}}+1\right)$ we obtain
$\sum_{j=0}^{N-1}\ln\left(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}+1\right)$ $=\int_{0}^{N-1}\ln(\,\mathrm{e}^{\alpha\cos\frac{2\pi}{N}x}+1)\mathop{}\!\mathrm{d}x+\frac{1}{2}[\ln(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{{\alpha}}+1)+\ln(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{{\alpha\cos\frac{2\pi}{N}}}+1)]+R$ (152)
$\approx\frac{N}{2\pi}\int_{0}^{2\pi-\frac{2\pi}{N}}\ln(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos x}+1)\mathop{}\!\mathrm{d}x+\ln(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{{\alpha}}+1)+\mathcal{O}(N^{-2})+R$ (153)
$=\frac{N}{2\pi}\left(\int_{0}^{2\pi}-\int_{2\pi-\frac{2\pi}{N}}^{2\pi}\right)\mathop{}\!\mathrm{d}x\,\ln(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos x}+1)+\ln(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{{\alpha}}+1)+R+\mathcal{O}(N^{-2})$ (154)
$=\frac{N}{2\pi}\int_{0}^{2\pi}\mathop{}\!\mathrm{d}x\,\ln(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos x}+1)+R+\mathcal{O}(N^{-2}).$ (155)
Now we estimate the remainder term $R$. We observe that (with $y=\frac{2\pi}{N}x$)
$\begin{split}\left\lvert f^{\prime}(x)\right\rvert&=\frac{2\pi\alpha}{N}\left\lvert\frac{\,\mathrm{e}^{\alpha\cos y}}{\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos y}+1}\sin y\right\rvert\\ &\leq\frac{2\pi\alpha}{N}\frac{\,\mathrm{e}^{\alpha\cos y}}{\left\lvert\,\mathrm{e}^{\alpha\cos y}+\,\mathrm{e}^{-\mathrm{i}\theta}\right\rvert}.\end{split}$ (156)
When the imaginary part of $\,\mathrm{e}^{\mathrm{i}\theta}$ is nonzero, i.e. $\sin\theta\neq 0$, then the denominator is lower-bounded by $\left\lvert\sin\theta\right\rvert$ and the numerator is upper-bounded by $\,\mathrm{e}^{\alpha}$, which imply
$|R_{0}|=\mathcal{O}(N^{-1}).$ (157)
Let us now consider $\sin\theta=0$, i.e. $\theta=0$ or $\pi$. For $\theta=0$, we have
$\begin{split}\left\lvert f^{\prime}(x)\right\rvert&\leq\frac{2\pi\alpha}{N}\frac{\,\mathrm{e}^{\alpha\cos y}}{{\,\mathrm{e}^{\alpha\cos y}+1}}\leq\frac{2\pi\alpha}{N}.\end{split}$ (158)
Therefore Eq. (157) still holds. Thus for any $\theta\neq\pi$ mod ${2\pi}$ we have shown that
$\sum_{j=0}^{N-1}\ln\left(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}+1\right)=\frac{N}{2\pi}\int_{0}^{2\pi}\mathop{}\!\mathrm{d}x\,\ln(\,\mathrm{e}^{\mathrm{i}\theta}\,\mathrm{e}^{\alpha\cos x}+1)+\mathcal{O}(N^{-1}).$ (159)
It is straightforward to extend the proof to the case where the sum is over half-integer values.
Next we deal with the special $\theta=\pi$ case, where the derivative $f^{\prime}(x)$ diverges at $x=\pi/2$. In fact, the same is true for any $f^{(2m+1)}(x)$ at $x=\pi/2$, so the formula is not particularly useful for any finite $m$. Thus we formally extend the formula to $m=\infty$:
$R_{\infty}\equiv R_{m=\infty}=\sum_{k=1}^{\infty}\frac{B_{2k}}{(2k)!}[f^{(2k-1)}(b)-f^{(2k-1)}(a)].$ (160)
It should be understood as an asymptotic series expansion of the remainder term. To get a finite result, we will need to resum the series.
[…]
$\sum_{j=\frac{1}{2}}^{\frac{N}{4}-\frac{1}{2}}\ln(\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}-1).$ (161)
Let $f(x)=\ln(\,\mathrm{e}^{\alpha\cos\frac{2\pi}{N}x}-1)$, then we have
$\sum_{j=\frac{1}{2}}^{\frac{N}{4}-\frac{1}{2}}\ln(\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}-1)$ $=\int_{\frac{1}{2}}^{\frac{N}{4}-\frac{1}{2}}\ln(\,\mathrm{e}^{\alpha\cos\frac{2\pi}{N}x}-1)\mathop{}\!\mathrm{d}x+\frac{1}{2}[\ln(\,\mathrm{e}^{{\alpha\cos\frac{\pi}{N}}}-1)+\ln(\,\mathrm{e}^{{\alpha\sin\frac{\pi}{N}}}-1)]+R_{\infty}$ (162)
$=\frac{N}{2\pi}\int_{\frac{\pi}{N}}^{\frac{\pi}{2}-\frac{\pi}{N}}\ln(\,\mathrm{e}^{\alpha\cos x}-1)\mathop{}\!\mathrm{d}x+\frac{1}{2}\Big{[}\ln(\,\mathrm{e}^{{\alpha}}-1)+\ln\frac{\alpha\pi}{N}\Big{]}+\mathcal{O}(N^{-2})+R_{\infty}$ (163)
$=\frac{N}{2\pi}\int_{0}^{\frac{\pi}{2}}\ln(\,\mathrm{e}^{\alpha\cos x}-1)\mathop{}\!\mathrm{d}x+\frac{1}{2}+R_{\infty}+\mathcal{O}(N^{-2}).$ (164)
Now we consider the remainder term. It is not difficult to show that
$f^{(2k-1)}\left(\frac{N}{4}-\frac{1}{2}\right)=\left(\frac{2\pi}{N}\right)^{2k-1}\frac{(2k-2)!}{(-\pi/N)^{2k-1}}=-2^{2k-1}(2k-2)!.$ (165)
And $f^{(2k-1)}(\frac{1}{2})$ is $\mathcal{O}(N^{-2k-2})$, so it can be ignored. Therefore
$R_{\infty}=-\sum_{k=1}^{\infty}\frac{B_{2k}}{2k(2k-1)}2^{2k-1}.$ (166)
Here we can see that the correction does not depend on $\alpha$, neither does it depend on the specific form of the spectrum considered. This explains why CFT results agree with results derived from lattice models at any finite temperature.
To resum the asymtotic series, we use the following integral representation of Bernouli numbers:
$B_{2k}=4k(-1)^{k+1}\int_{0}^{\infty}\frac{t^{2k-1}}{\,\mathrm{e}^{2\pi t}-1}\mathop{}\!\mathrm{d}t.$ (167)
Then we find
$R_{\infty}=-2\int_{0}^{\infty}\mathop{}\!\mathrm{d}t\,\frac{1}{\,\mathrm{e}^{2\pi t}-1}\sum_{k=1}^{\infty}(-1)^{k+1}\frac{(2t)^{2k-1}}{2k-1}=-2\int_{0}^{\infty}\mathop{}\!\mathrm{d}t\,\frac{\arctan 2t}{\,\mathrm{e}^{2\pi t}-1}=-\left(\frac{1}{2}-\ln\sqrt{2}\right).$ (168)
The last integral is given by Binet’s second formula for the logarithm of the Gamma function Whittaker and Watson (1996):
$2\int_{0}^{\infty}\mathop{}\!\mathrm{d}t\,\frac{1}{\,\mathrm{e}^{2\pi t}-1}\arctan\frac{t}{z}=z-\left(z-\frac{1}{2}\right)\ln z+\ln\Gamma\left(z\right)-\ln\sqrt{2\pi}.$ (169)
Setting $z=\frac{1}{2}$ gives the desired result.
So putting together we have shown that
$\sum_{j=\frac{1}{2}}^{\frac{N}{4}-\frac{1}{2}}\ln(\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}-1)=\frac{N}{2\pi}\int_{0}^{\frac{\pi}{2}}\mathop{}\!\mathrm{d}x\,\ln(\,\mathrm{e}^{\alpha\cos x}-1)+\ln\sqrt{2}+\mathcal{O}(N^{-2}).$ (170)
Similarly,
$\begin{split}\sum_{j=\frac{N}{4}+\frac{1}{2}}^{\frac{N}{2}-\frac{1}{2}}\ln(1-\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}})&=\sum_{j=\frac{N}{4}+\frac{1}{2}}^{\frac{N}{2}-\frac{1}{2}}\ln\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}(\,\mathrm{e}^{-\alpha\cos\frac{2\pi j}{N}}-1)\\ &=\alpha\sum_{j=\frac{N}{4}+\frac{1}{2}}^{\frac{N}{2}-\frac{1}{2}}\cos\frac{2\pi j}{N}+\sum_{j=\frac{N}{4}+\frac{1}{2}}^{\frac{N}{2}-\frac{1}{2}}\ln(\,\mathrm{e}^{-\alpha\cos\frac{2\pi j}{N}}-1)\\ &=-\frac{\alpha}{2\sin\frac{\pi}{N}}+\sum_{j=\frac{1}{2}}^{\frac{N}{4}-\frac{1}{2}}\ln(\,\mathrm{e}^{-\alpha\cos\frac{2\pi j}{N}}-1)\\ &=-\frac{\alpha}{2\sin\frac{\pi}{N}}+\frac{N}{2\pi}\int_{0}^{\frac{\pi}{2}}\mathop{}\!\mathrm{d}x\,\ln(\,\mathrm{e}^{\alpha\cos x}-1)+\ln\sqrt{2}+\mathcal{O}(N^{-2})\\ &=\frac{N}{2\pi}\int_{\frac{\pi}{2}}^{\pi}\mathop{}\!\mathrm{d}x\,\ln(1-\,\mathrm{e}^{\alpha\cos x})+\ln\sqrt{2}+\mathcal{O}(N^{-2}).\end{split}$ (171)
A straightforward generalization is to consider $f(x)=\ln\left(\,\mathrm{e}^{\alpha\cos\frac{2\pi}{N}x-\alpha\cos\left(\pi\nu\right)}-1\right)$ and the summation
$\sum_{j=\frac{1}{2}s^{\prime}}^{N-1+\frac{1}{2}s^{\prime}}f(j),$ (172)
where $\nu\in(0,1)$ and $s^{\prime}\in\left\{0,1\right\}$. Now the singular points are at $x_{1}=\frac{N\nu}{2}$ and $x_{2}=N-\frac{N\nu}{2}$. Assume the largest $j<x_{1}$ in the summation is $j_{1}=x_{1}-\delta$. Then, we have
$\sum_{j=\frac{1}{2}s^{\prime}}^{x_{1}-\delta}f(j)$ $=\int_{\frac{1}{2}s^{\prime}}^{x_{1}-\delta}f(x)\mathop{}\!\mathrm{d}x+\frac{1}{2}\left[f(\frac{1}{2}s^{\prime})+f(x_{1}-\delta)\right]+R_{\infty}+\mathcal{O}(N^{-2})$ (173)
$=\int_{0}^{x_{1}}f(x)\mathop{}\!\mathrm{d}x+\left(\frac{1}{2}-\delta\right)\ln\left[\alpha\sin(\pi\nu)\frac{2\pi\delta}{N}\right]+\delta+R_{\infty}+\mathcal{O}(N^{-2}).$ (174)
The remainder term $R_{\infty}$ is
$R_{\infty}=-\sum_{k=1}^{\infty}\frac{B_{2k}}{2k(2k-1)}\left(\frac{1}{\delta}\right)^{2k-1}=-\delta+\left(\delta-\frac{1}{2}\right)\ln\delta+\ln\frac{\sqrt{2\pi}}{\Gamma(\delta)}.$ (175)
Thus,
$\sum_{j=\frac{1}{2}s^{\prime}}^{x_{1}-\delta}f(j)=\int_{0}^{x_{1}}f(x)\mathop{}\!\mathrm{d}x+\left(\frac{1}{2}-\delta\right)\ln\left[\alpha\sin(\pi\nu)\frac{2\pi\delta}{N}\right]+\left(\delta-\frac{1}{2}\right)\ln\delta+\ln\frac{\sqrt{2\pi}}{\Gamma(\delta)}.$ (176)
If we add up the constant corrections from four segments, we will find
$\sum_{j=\frac{1}{2}s^{\prime}}^{N-1+\frac{1}{2}s^{\prime}}\ln\left\lvert\,\mathrm{e}^{\alpha\cos\frac{2\pi}{N}j-\alpha\cos\left(\pi\nu\right)}-1\right\rvert=\frac{N}{2\pi}\int_{0}^{2\pi}\ln\left\lvert\,\mathrm{e}^{\alpha\cos x-\alpha\cos\left(\pi\nu\right)}-1\right\rvert\mathop{}\!\mathrm{d}x+2\ln\left[2\sin(\pi\delta)\right].$ (177)